OpenStudy (babynini):

a) sketch the curve b) find a rectangular coordinate for the curve by eliminating the perimeter x=tant, y=cot t 0

OpenStudy (babynini):

i did some stuff but think I just complicated it more:P

OpenStudy (babynini):

@sdfgsdfgs

OpenStudy (babynini):

@IrishBoy123

OpenStudy (babynini):

by ^2 both sides of x and adding 1 I got 1+x^2=sec^2t But i'm not sure if that made it better or worse

OpenStudy (irishboy123):

sketch what curve??

OpenStudy (babynini):

The curve we'll get from the equation we make. It says "sketch the curve represented by the parametric equations" but I find it's easier to start with b

OpenStudy (babynini):

@IrishBoy123

OpenStudy (babynini):

@dan815 :)

OpenStudy (irishboy123):

so "x=tant, y=cot t " you see anything in that?!?!

OpenStudy (babynini):

er well it's the reciprocal

OpenStudy (babynini):

x=sint/cost y=1/(sint/cost)

OpenStudy (babynini):

oo so y=1/x ??

OpenStudy (babynini):

@IrishBoy123

OpenStudy (sdfgsdfgs):

yes @Babynini :)

OpenStudy (babynini):

Thanks :)